Find the inverse equation of f(x)=(x^3-7)^5.
Answer by CubeyThePenguin(3113)

Solve for x.

y = (x^3 - 7)^5
y^5 = x^3 - 7
y^5 + 7 = x^3
cbrt(y^5 + 7) = x

f^(-1)(x) = cbrt(x^5 + 7)